# Double Configuration Interaction Singles: Scalable and size-intensive approach for orbital relaxation in excited states and bond-dissociation

### v1

@@ III.3 Single bond dissociation
Due to the inclusion of both HF and doubly excited configurations in DCIS, it is expected to provide reasonable descriptions for single bond breaking, which is challenging for many existing methods. This is indeed the case: Fig. 3 presents the potential energy curves for (a) hydrogen fluoride and (b) diatomic fluorine molecules computed with a 6-31G basis, along with the energy errors from the full CI (FCI) shown in (c) and (d). DCIS successfully yields curves that are parallel to FCI, with non-parallelity errors (NPEs) of 14.7 mH and 15.9 mH for HF and F2, respectively. Notably, the DCIS ground state energy is consistently lower than that of HF, but there is limited correlation effect, indicating that the method is size-intensive and primarily captures static correlation. The de-excitation of a CIS state simply reverts to the HF state when static correlation is absent. Importantly, these results suggest that CIS states can serve as a reasonably accurate reference even when the ground state HF state fails at the dissociation limit.
Fig. 3 also depicts the results of SFCIS, another CIS variant known for enabling single-bond breaking in a size-intensive manner. Since SFCIS employs a high-spin triplet state as a reference, the orbitals are suitable at dissociation limit but not for the singlet ground state around equilibrium. Hence, the SFCIS energy tends to be relatively lower at the former compared to the latter, resulting in the considerably shallow potential curves (NPEs are 81.7 mH and 38.4 mH, see Fig. 3).
@@ IV Concluding remarks
In this work, we first derived the CIS electronic Hessian, and building upon this foundation, introduced a new size-intensive method, Double CIS. DCIS variationally handles the orbital relaxation of the target CIS state at a mean-field cost. Furthermore, we proposed an efficient approach for the direct optimization in DCIS, particularly useful for high-lying excited states, by evading the determination of lower-lying states. Our results show significant improvements in the descriptions of charge transfer excitations and bond breaking systems. We anticipate that incorporating dynamical correlation will further enhance these results, which we plan to address in future work.
@@ Acknowledgement
This work was supported by the start-up funding at Shibaura Institute of Technology and JST FOREST Program, Grant No. JPMJFR223U.

### v3

@@ III.3 Single bond dissociation
Due to the inclusion of both HF and doubly excited configurations in DCIS, it is expected to provide reasonable descriptions of single bond breaking, which is challenging for many existing methods. This is indeed the case: Fig. 5 presents the potential energy curves for (a) hydrogen fluoride and (b) diatomic fluorine molecules computed with a 6-31G basis, along with the energy errors from full CI (FCI) shown in (c) and (d). DCIS successfully yields reasonable potential energy curves. Notably, the DCIS ground state energy is consistently lower than that of HF, but its correlation effect is limited near the equilibrium bond length, highlighting that the method is size-intensive and primarily captures static correlation. The de-excitation of a CIS state simply reverts to the HF state when static correlation is absent. Importantly, these findings suggest that CIS states can serve as reasonably accurate references even when the ground state HF state fails at the dissociation limit.
In Table 3, the non-parallelity errors (NPEs) against FCI potential curves are summarized. Interestingly, DCIS achieves NPEs of 14.7 mH and 15.9 mH for HF and F2, respectively, which are comparable to—or even smaller than—those of standard CASSCF with a 2-electron-2-orbital active space (20.3 and 24.6 mH). However, we consider this may result from error cancellation in DCIS. While CASSCF optimizes orbitals self-consistently and generally yields shallow potential curves, DCIS lacks orbital optimization, resulting in higher energy than CASSCF at the dissociation limit. Overall, DCIS cancels out the effects of dynamical correlation and orbital optimization, leading to relatively balanced NPEs.
Fig. 5 also depicts the results of SFCIS, another CIS variant known for enabling single-bond breaking in a size-intensive manner. Since SFCIS employs a high-spin triplet state as a reference, the orbitals are well-suited for the dissociation limit but not necessarily for the singlet ground state near equilibrium. Consequently, SFCIS energies tend to be relatively lower at dissociation, resulting in considerably shallow potential curves. The corresponding NPEs are 81.7 mH and 38.4 mH.
@@
III.4 Convergence behavior: maximum-overlap v.s. block Davidson
Many of the excited states discussed in Section III.2 are high-lying, and can be accessed from the bottom using the Davidson algorithm in CIS. As described in Section II.2.2, the DCIS solutions also generally appear as higher roots, presenting a similar challenge.
In this section, we examine the convergence behavior of DCIS using the maximum-overlap Davidson algorithm, comparing it to the block Davidson method across various types of excitations. For this analysis, we selected three systems from the test set used above: one excited state of benzonitrile (4) and two excited states of $\beta$-Dipeptide (9). The third CIS excited state of benzonitrile is relatively valence in character, with the CIS orbital gradient being $|^{\rm o}{\bf g}|=0.106$ and a dipole change of $|\Delta{\bm{\mu}}|=1.01$. The two excited states of $\beta$-Dipeptide corresponding to the 13th and 46th CIS eigenstates both exhibit large gradients and dipole changes, $|^{\rm o}{\bf g}|=0.257$ and $0.290$ and $|\Delta{\bm{\mu}}|=2.60$ and $4.65$, respectively.
In the maximum-overlap Davidson algorithm, the initial trial vector is set to the CI coefficients of the target CIS state $I$, i.e., $\bar{\bf d}={\bf c}^{I}$ and ${\bf d}={\bf 0}$. For the block Davidson algorithm, $(I+1)$ initial states (including the HF state) are prepared, with trial vectors chosen as the CI coefficients of the target state and all lower-lying CIS states. The convergence criterion is set to ensure the DCIS gradient norm is smaller than $10^{-5}$.
The results are shown in Fig. 6, which plots the gradient norm of each solution as a function of the number of trial vectors (i.e., subspace dimension). The block Davidson algorithm handles multiple solutions at each iteration, and often provides stable convergence for all states simultaneously. In practice, the number of cycles required to converge all states is 10-15, regardless of the system. However, a major drawback is the rapid expansion of the subspace; namely, at each iteration, the dimension increases by $I+1$ additional trial vectors. As a result, the computational cost required to converge the target state becomes increasingly larger for higher-lying CIS states. For example, while the $S_{3}$ state of 4 (Fig. 6(a)) requires a subspace with 41 vectors to converge the target state (red curve), the $S_{46}$ state of 9 (Fig. 6(c)) requires 396 vectors (the number of cycles before convergence is 13 and 10, respectively).
In contrast, it is clearly shown in Fig. 6 that the maximum-overlap algorithm (green curve) converges with a much smaller subspace dimension compared to the block Davidson algorithm. For the $S_{3}$ state of 4, it takes only 14 cycles (vectors), achieving a threefold cost reduction compared to the block Davidson method. However, when the target state has a larger $|^{\rm o}{\bf g}|$, as in the two excitations of 9, it generally requires more iterations and vectors to converge, due to the larger relaxation effect. Specifically, the maximum-overlap method converges with 30 and 45 vectors for $S_{13}$ and $S_{46}$ of (9), respectively. At some iterations, the DCIS gradient increases compared to the previous iteration, typically indicating the emergence of intruder (low-lying) state in the subspace. Nonetheless, convergence is still achieved stably, and the sizes of these subspaces are 4.2 and 8.8 times smaller than those of the block Davidson method, offering a significant reduction in cost.
@@ IV Concluding remarks
In this work, we first derived the CIS electronic Hessian, and building upon this foundation, introduced a new size-intensive method, Double CIS. DCIS approximately handles the orbital relaxation of the target CIS state to first order. Thanks to the variational nature of DCIS, it can be expressed as a generalized eigenvalue problem, effectively representing “CIS on top of CIS,” which enables stable optimization in contrast to the challenges associated with the non-linear full orbital optimization. Furthermore, we proposed an efficient approach for the direct optimization in DCIS, particularly useful for high-lying excited states, by evading the determination of lower-lying states. Our results show significant improvements in the descriptions of charge transfer excitations and bond breaking systems. We anticipate that incorporating dynamical correlation will further enhance these results, which we plan to address in future work.
@@ Acknowledgement
This work was supported by the start-up funding from the Shibaura Institute of Technology and JST FOREST Program, Grant No. JPMJFR223U.
@@
Supplementary Material
The Supplementary Material includes: (1) a detailed derivation of the analytical expression for the CIS Hessian, (2) a discussion on the equivalence of different DCIS parameterizations, (3) the pseudocode for the maximum-overlap Davidson algorithm, (4) a table listing excitation energies, and (5) tables presenting the total energies of the HF and F2 molecules.